# intersection.bib

### v2

@@ 1 Introduction
Characteristics of road networks have been considered in numerous works. As in this paper we are interested in a mathematical model to construct road networks, we only mention two papers, first Barthélemy and Flammini [bar_fla_mod_urban_st_patt], who build urban networks by adding new centres and connecting them to the existing network in certain optimal way. Second, Molinero and Hernando [mol_her_model_4_gen_rd] introduce an algorithm based on a succession of Delanuay triangulation of a dense network of the centres of interest plus auxiliary points, and only keeping paths that are shortest between the original nodes of interest in the resulting network.
@@
One of the fundamental objects in FPP are geodesics, paths that collect minimal costs connecting two points among a random penalty landscape on the plane. These will be our candidates to model roads. For quantitative analysis of our road model we need refined estimates on geodesics, which unfortunately are not yet available for FPP models. Instead, we switch to exponential last passage percolation (LPP), where extensive results on last passage times and geometry of geodesics are now available. FPP is widely believed to belong to the KPZ universality class – a property proved for LPP. This implies that the geometry of geodesics in the two models should share basic characteristics that we use throughout our arguments. Thus, switching to LPP from FPP seems a reasonable move. Of the vast literature on the KPZ universality class, we refer to the surveys by Corwin [corwin_kpz_univ16], Ferrari-Spohn [fer-spohn_rnd_growth11], and Quastel [quas_intro_kpz12].
@@ 1.1 A Poisson model for road network
Without loss of generality we now assume that a car starts from coordinate $(-z,0)$ of $\mathbb{R}^{2}$, where $z$ is a positive real. If $z\leq r$ then the car does appear in the disc $B(0,\,r)$ of radius $r$ around the origin.
@@
The mean distance travelled by cars is $\frac{1}{\gamma}$, and since we already fixed the Poisson intensity of cars at 1, we can think of this as a large number. Similarly, the radius $r$ of interest e.g., for a homeowner wishing for a quiet house, can be considered $\mathcal{O}(1)$. Hence we find that even in a moderate sized garden, plenty of cars should pass. (For actual figures, the UK’s 243,610 $\text{km}^{2}$ and 33.5m cars give 85 metres as unit length to get density 1 of cars, and the typical driving distance is around 10…15 km: $\frac{1}{\gamma}$ is in the order of 100’s. Of course spatial density fluctuates drasticly between urban and rural areas.) Taking the driving distance to infinity ($\gamma\searrow 0$) makes $N_{r}$ divergent.
@@ 2.1 Further notations and results
Let $\epsilon>0$ be fixed; from now on all the constants we will obtain from the theorems will depend only on this $\epsilon$. From each vertex $v\in\mathbb{Z}^{2}$ an angle is chosen according to i.i.d. uniform $(\epsilon,\frac{\pi}{2}-\epsilon)$ distribution and we consider the a.s. unique semi-infinite geodesic in the chosen direction. We will denote the angle chosen by $v$ by $\theta_{v}$ and the semi-infinite geodesic in that direction by $\Gamma_{v}^{\theta_{v}}$. For $r\in\mathbb{Z}$ let $\boldsymbol{r}$ denote the vertex $(r,r)\in\mathbb{Z}^{2}$. For convenience sometimes we will work with the rotated axes $x+y=0$ and $x-y=0$. They will be called space axis and time axis respectively. For a vertex $v\in\mathbb{Z}^{2}$, $\phi(v)$ will denote the time coordinate of $v$ and $\psi(v)$ will denote the space coordinate of $v$. Precisely, for $u,v\in\mathbb{Z}^{2}$, we have
@@
Let $N$ denote the number of cars passing through origin. More precisely we define,
$N:=|\{u\in\mathbb{Z}^{2}:\boldsymbol{0}\in\Gamma_{u}^{\theta_{u}}\}|.$
@@
But $N$ has infinite expectation. To show this we first need the following definition. For $n\in\mathbb{N}$, define $N_{n}$ to be the number of cars starting from the line $\mathcal{L}_{-n}$ and going through origin. We have the following theorem.
@@
For $x\in\mathcal{L}_{-n},$ let $\mathcal{G}_{x}$ denote the event that the car starting from $x$ goes through origin. Then
@@
There exist constants $C_{1},c_{1}>0$ (depending on $\epsilon$) and $n$ sufficiently large such that
@@
There exists $C,c>0$ (depending on $\epsilon$) such that the following holds
@@
For $\delta>0$ sufficiently small (depending on $\epsilon$) we have there exists constant $C>0$ (depending on $\epsilon$) and for sufficiently large $n$ (depending on $\epsilon$)
$\mathbb{P}(T_{n}\geq\frac{\delta n^{1/3}}{\log n})\geq 1-C\delta.$
@@
The proof follows from a union bound. let $V$ denote the line segments with $\lceil{\frac{\delta n^{1/3}}{\log n}}\rceil$ vertices and with midpoint $\boldsymbol{0}$ on the line $\mathcal{L}_{0}$. Then
$\{T_{n}\leq\frac{\delta n^{1/3}}{\log n}\}\subset\bigcup_{v\in V}\{N_{v}\geq n^{4/3}\}.$
Hence, by Theorem 2.4 upper bound we have for some constant $C>0$ (depending on $\epsilon$)
$\mathbb{P}(T_{n}\leq\frac{\delta n^{1/3}}{\log n})\leq C\delta.$
@@
The above theorem says that, with high probability, to see a vertex having $n^{4/3}$ many cars through it, we need to go at least sufficiently small multiple of $\frac{n^{1/3}}{\log n}$ distance from origin. The following theorem shows that, with non-vanishing probability, there is at least one vertex within $n^{1/3}$ distance around the origin that has $n^{4/3}$ many cars through it. Precisely, for a fixed constant $\ell_{0}$ we make a slight modification to the definition of $T_{n}$ and by abuse of notation we will still call the new variable $T_{n}.$
$T_{n}:=\min\{|\psi(v)|:v\in\mathcal{L}_{0},N_{v}\geq\frac{n^{4/3}}{\ell_{0}}\}.$
@@
There exists $c>0,\ell_{0}$ (depending on $\epsilon$) such that for sufficiently large $n$ (depending on $\epsilon$) we have
$\mathbb{P}(T_{n}\leq n^{1/3})\geq c.$
@@ 3 Bounds for the furthest distance a car can come from
First we divide the interval $(\epsilon,\frac{\pi}{2}-\epsilon)$ into disjoint sub-intervals. Precisely, for $1\leq i\leq\lfloor n^{1/3}\rfloor$ and for a uniformly bounded sequence $\{c_{n}\}$ we partition the interval $(\epsilon,\frac{\pi}{2}-\epsilon$) into intervals $A_{i}$ each of equal length $\frac{c_{n}(\frac{\pi}{2}-2\epsilon)}{n^{1/3}}$. For $1\leq i\leq\lfloor n^{1/3}\rfloor$ we define the following random variables.
$D_{i}:=\max\{n\in\mathbb{N}:\exists u\in\mathcal{L}_{-n}\text{ such that }\theta_{u}\in A_{i}\text{ and }\boldsymbol{0}\in\Gamma_{u}^{\theta_{u}}\}.$
@@
$D_{i}^{v}:=\max\{n\in\mathbb{N}:\exists u\in\mathcal{L}_{-n}\text{ such that }\theta_{u}\in A_{i}\text{ and }v\in\Gamma_{u}^{\theta_{u}}\}.$
Clearly, for all $v\in V$ we have
@@
Figure 3: To prove the upper bound in Theorem 2.5 we fix $1\leq i\leq\lfloor n^{1/3}\rfloor$. The event that there are more than $\ell$ many $u\in\mathcal{L}_{-n}$ with $\theta_{u}\in A_{i}$ and $\Gamma_{u}^{\theta_{u}}$ intersecting $V$ can happen in two ways. Either the above $u$ lies in $\mathcal{L}_{-n}\setminus\widetilde{V}$. For sufficiently large $\ell$ this will imply large transversal fluctuation for either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$. Proposition 3.3 gives a stretched exponential upper bound for this event. The other possibility is there are more than $\ell$ distinct vertices inside on $V$ that carries a geodesic from $\widetilde{V}$. Proposition 3.4 gives a stretched exponential upper bound for this event.
@@
Now, by planarity the event $\mathcal{A}_{i}$ implies either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$ will intersect $V$ (see Figure 3). Since, $(\beta_{i}-\alpha_{i})$ is of order $\frac{1}{n^{1/3}}$, for sufficiently large $\ell$ (depending on $\epsilon$), the event $\mathcal{A}_{i}$ will imply that either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$ will have transversal fluctuation larger than $\frac{\ell^{1/96}n^{2/3}}{4}$. We have the following proposition.
Proposition 3.3.
[BBB23, Proposition 1.5] For $\epsilon>0$ there exist $C_{1},c_{1}>0$ such that for all $T>0,\ell>0,n\geq 1$ and for all $\alpha\in(\epsilon,\frac{\pi}{2}-\epsilon)$ we have
  1. (i)
  2. (ii)
Hence, for sufficiently large $\ell$ we have
@@
Now, we consider the event $\{\widehat{D_{i}}\geq\ell\}\cap\mathcal{A}_{i}^{c}$. On this event there are at least $\ell$ distinct vertices on $V$ that has a geodesic coming from $\widetilde{V}.$ We have the following proposition.
Proposition 3.4.
$\mathbb{P}(M_{n}^{k}\geq\ell)\leq Ce^{-c\ell^{1/128}}.$ (3.5)
Hence, using Proposition 3.4 we have for sufficiently large $\ell$ and $n$
@@
So, finally we have there exists $\ell_{0}$ (depending only on $\epsilon$) and $\widetilde{C}$ (depending only on $\epsilon$) such that for all large $n$
@@
So, from (3.2) we have there exists $C>0$ (depending only on $\epsilon$) such that for all $1\leq i\leq\lfloor n^{1/3}\rfloor$ we have
@@
We again consider disjoint sub-intervals of $(\epsilon,\frac{\pi}{2}-\epsilon)$. First we fix large constant $M$ which will be chosen later. For $1\leq i\leq[\frac{n^{1/3}}{2M}]$, we consider the intervals $A_{i}\subset(\epsilon,\frac{\pi}{2}-\epsilon)$, each of length $\frac{(\frac{\pi}{2}-2\epsilon)}{n^{1/3}}$ and midpoint $\frac{Mi(\frac{\pi}{2}-2\epsilon)}{n^{1/3}}$. Note that, unlike the previous case we are not partitioning the interval $(\epsilon,\frac{\pi}{2}-\epsilon$) this time. Instead we are taking points which are $\frac{M(\frac{\pi}{2}-2\epsilon)}{n^{1/3}}$ distance apart from each other and we are taking intervals of length $\frac{(\frac{\pi}{2}-2\epsilon)}{n^{1/3}}$ around them. We do this so that the constants that we get at the end depend only on $\epsilon$ but not on $M$. This will be elaborated later. Corresponding to each $A_{i}$ we define the following random variables.
$D_{i}:=\max\{n\in\mathbb{N}:\exists u\in\mathcal{L}_{-n}\text{ such that }\theta_{u}\in A_{i}\text{ and }\boldsymbol{0}\in\Gamma_{u}^{\theta_{u}}\}.$
Clearly,
@@
For $1\leq i\leq[\frac{n^{1/3}}{2M}]$ let
$\mathcal{B}_{i}:=\{D_{i}\geq n\}.$
So, applying inclusion-exclusion principle and considering the first two terms we have
$\mathbb{P}(D\geq n)\geq\sum_{i=1}^{[\frac{n^{1/3}}{2M}]}\mathbb{P}(\mathcal{B}_{i})-\sum_{1\leq i<j\leq[\frac{n^{1/3}}{2M}]}\mathbb{P}(\mathcal{B}_{i}\cap\mathcal{B}_{j}).$ (3.7)
@@
$D_{i}^{v}:=\max\{n\in\mathbb{N}:\exists u\in\mathcal{L}_{-n}\text{ such that }\theta_{u}\in A_{i}\text{ and }v\in\Gamma_{u}^{\theta_{u}}\}.$
Figure 4: To prove the lower bound for the first sum in (3.7) we fix an $i$. Using Proposition 3.3 we can chose $M$ large enough so that on a positive probability event $\Gamma_{v_{1}^{\prime}}^{\beta_{i}}$ and $\Gamma_{v_{2}^{\prime}}^{\alpha_{i}}$ intersect $\mathcal{L}_{0}$ on $V_{M}$. Further, we consider an independent event on which there is at least one vertex on $V^{\prime}$ that chooses an angle in $A_{i}$. Due to central limit theorem the later event is a positive probability event. Hence, the intersection of these two events is a positive probability event and on this intersection $\widehat{D}_{i}\geq 1$.
@@
$\mathcal{C}_{i}:=\{\Gamma_{v_{1}^{\prime}}^{\beta_{i}}(0)\in V_{M}\text{ and }\Gamma_{v_{2}^{\prime}}^{\alpha_{i}}(0)\in V_{M}\}.$
Again as $(\beta_{i}-\alpha_{i})$ is of order $\frac{1}{n^{1/3}},$ the complement of the event $\mathcal{C}_{i}$ will imply that either $\Gamma_{v_{1}^{\prime}}^{\beta_{i}}$ or $\Gamma_{v_{2}^{\prime}}^{\alpha_{i}}$ will have transversal fluctuation more than $\frac{M}{2}n^{2/3}$. So, using Proposition 3.3 we can chose $M$ large enough so that
@@
$\mathcal{D}_{i}:=\{\exists u\in V^{\prime}\text{ such that }\theta_{u}\in A_{i}\}=\{\sum_{u\in V^{\prime}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}\geq 1\}.$
So, we consider the random variable $X_{i}:=\sum_{u\in V^{\prime}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}.$ We have $\mathbb{E}(X_{n})=\frac{c_{n}\lceil n^{2/3}\rceil}{n^{1/3}}$ and $\operatorname{{\mathbb{V}ar}}(X_{n})=(\frac{c_{n}}{n^{1/3}}-\frac{c_{n}^{2}}{n^{2/3}})\lceil n^{2/3}\rceil.$ By the Berry-Esseen inequality we have for all $n$ there exists a constant $c>0$ such that
$|\mathbb{P}(X_{n}-\mathbb{E}(X_{n})\geq 0\})-\frac{1}{2}|\leq\frac{c}{(c_{n}n^{1/3}-c_{n}^{2})n^{1/3}}.$
Hence, for sufficiently large $n$ we have
$\mathbb{P}(X_{n}-\mathbb{E}(X_{n})\geq 0\})\geq\frac{1}{4}.$
@@
Clearly, by planarity of geodesics (see Section 2 for definition), on the event $\mathcal{C}_{i}\cap\mathcal{D}_{i},\widehat{D_{i}}\geq 1.$ So, from (3.8) we have there exists $c_{1}>0$ such that

### v3

@@ 1 Introduction
Characteristics of road networks have been considered in numerous works. As in this paper we are interested in a mathematical model to construct road networks, we only mention a handful of papers, first Barthélemy and Flammini [bar_fla_mod_urban_st_patt], who build urban networks by adding new centres and connecting them to the existing network in certain optimal way.
Aldous [aldous_scaleinv_sp_nw] introduced the notion of Scale-invariant random spatial networks as a potential model for road networks. This was followed by a natural construction of such an object by Kendall [kendall_rnd_lines_ms] and by Kahn [kahn_impr_poi_lines] based on a Poisson line process. We will base our model on a directed KPZ-style construction, hence we do not expect it to be scale-invariant.
Molinero and Hernando [mol_her_model_4_gen_rd] introduce an algorithm based on a succession of Delanuay triangulation of a dense network of the centres of interest plus auxiliary points, and only keeping paths that are shortest between the original nodes of interest in the resulting network.
@@
One of the fundamental objects in FPP are geodesics, paths that collect minimal costs connecting two points among a random penalty landscape on the plane. These will be our candidates to model roads. In fact [HW65] already asked about “highways and byways” in terms of edges frequented by geodesics, and the question received a partial answer by Ahlberg, Hanson and Hoffman [ahl_hans_hoff_no_geo_fpp].
For quantitative analysis of our road model we need refined estimates on geodesics, which unfortunately are not yet available for FPP models. Instead, we switch to exponential last passage percolation (LPP), where extensive results on last passage times and geometry of geodesics are now available. FPP is widely believed to belong to the KPZ universality class – a property proved for LPP. This implies that the geometry of geodesics in the two models should share basic characteristics that we use throughout our arguments. Thus, switching to LPP from FPP seems a reasonable move. Of the vast literature on the KPZ universality class, we refer to the surveys by Corwin [corwin_kpz_univ16], Ferrari-Spohn [fer-spohn_rnd_growth11], and Quastel [quas_intro_kpz12].
@@ 1.1 A Poisson model for road network
Without loss of generality we now assume that a car starts from coordinate $(-z,0)$ of $\mathbb{R}^{2}$, where $z$ is a positive real. If $z\leq r$ then the car already starts inside the disc $B(0,\,r)$ of radius $r$ around the origin.
@@
The mean distance travelled by cars is $\frac{1}{\gamma}$, and since we already fixed the Poisson intensity of cars at 1, we can think of this as a large number. (For actual figures, if the UK’s 33.5m cars formed a homogeneous Poisson process on the UK’s 243,610 $\text{km}^{2}$ area, then they would be spaced an average of 43 metres apart. The typical driving distance is around 10…15 km, which is far longer. Hence $\frac{1}{\gamma}$ can be assumed to be much greater than $\frac{1}{2}$, which is the average distance between Poisson points of intensity 1, using the units of this article. Of course spatial density fluctuates drasticly between urban and rural areas.) Similarly, the radius $r$ of interest e.g., for a homeowner wishing for a quiet house, can be considered $\mathcal{O}(1)$. Hence we find that even in a moderate sized garden, plenty of cars should pass. Taking the driving distance to infinity ($\gamma\searrow 0$) makes $N_{r}$ divergent.
@@ 2.1 Further notations and results
Let $\varepsilon>0$ be fixed; from now on all the constants we will obtain from the theorems will depend only on this $\varepsilon$. From each vertex $v\in\mathbb{Z}^{2}$ an angle is chosen according to i.i.d. uniform $(\varepsilon,\frac{\pi}{2}-\varepsilon)$ distribution and we consider the a.s. unique semi-infinite geodesic in the chosen direction. We will denote the angle chosen by $v$ by $\theta_{v}$ and the semi-infinite geodesic in that direction by $\Gamma_{v}^{\theta_{v}}$. For $r\in\mathbb{Z}$ let $\boldsymbol{r}$ denote the vertex $(r,r)\in\mathbb{Z}^{2}$. For convenience sometimes we will work with the rotated axes $x+y=0$ and $x-y=0$. They will be called space axis and time axis respectively. For a vertex $v\in\mathbb{Z}^{2}$, $\phi(v)$ will denote the time coordinate of $v$ and $\psi(v)$ will denote the space coordinate of $v$. Precisely, for $u,v\in\mathbb{Z}^{2}$, we have
@@
Let $N$ denote the number of cars passing through the origin. More precisely we define,
$N:=\#\{u\in\mathbb{Z}^{2}:\boldsymbol{0}\in\Gamma_{u}^{\theta_{u}}\}.$
@@
But $N$ has infinite expectation. To show this we first need the following definition. For $n\in\mathbb{N}$, define $N_{n}$ to be the number of cars starting from the line $\mathcal{L}_{-n}$ and going through the origin. We have the following theorem.
@@
For $x\in\mathcal{L}_{-n},$ let $\mathcal{G}_{x}$ denote the event that the car starting from $x$ goes through the origin. Then
@@
There exist constants $C_{1},c_{1}>0$ (depending on $\varepsilon$) and $n$ sufficiently large such that
@@
There exists $C,c>0$ (depending on $\varepsilon$) such that the following holds
@@
For $\delta>0$ sufficiently small (depending on $\varepsilon$) we have that there exists constant $C>0$ (depending on $\varepsilon$) and for sufficiently large $n$ (depending on $\varepsilon$)
$\mathbb{P}\left(T_{n}\geq\frac{\delta n^{1/3}}{\log n}\right)\geq 1-C\delta.$
In particular, this implies there exists $c$ (depending on $\varepsilon$) such that
$\mathbb{E}\left(T_{n}\right)\geq\frac{cn^{1/3}}{\log n}.$
@@
The proof follows from a union bound. Let $V$ denote the line segment with $\lceil{\frac{\delta n^{1/3}}{\log n}}\rceil$ many vertices and with midpoint $\boldsymbol{0}$ on the line $\mathcal{L}_{0}$. Then
$\left\{T_{n}\leq\frac{\delta n^{1/3}}{\log n}\right\}\subset\bigcup_{v\in V}\left\{N_{v}\geq n^{4/3}\right\}.$
Hence, by Theorem 2.4 upper bound we have for some constant $C>0$ (depending on $\varepsilon$)
$\mathbb{P}\left(T_{n}\leq\frac{\delta n^{1/3}}{\log n}\right)\leq C\delta.$
@@
The above theorem says that, with high probability, to see a vertex having $n^{4/3}$ many cars through it, we need to go at least sufficiently small multiple of $\frac{n^{1/3}}{\log n}$ distance from origin. The following theorem shows that, with non-vanishing probability, there is at least one vertex within $n^{1/3}$ distance around the origin that has $n^{4/3}$ many cars through it. Precisely, for a fixed constant $\ell_{0}$ we make a slight modification to the definition of $T_{n}$ and call the new variable $T^{\ell_{0}}_{n}$.
$T^{\ell_{0}}_{n}:=\min\left\{|\psi(v)|:v\in\mathcal{L}_{0},N_{v}\geq\frac{n^{4/3}}{\ell_{0}}\right\}.$
@@
There exists $c>0,\ell_{0}$ (depending on $\varepsilon$) such that for sufficiently large $n$ (depending on $\varepsilon$) we have
$\mathbb{P}(T^{\ell_{0}}_{n}\leq n^{1/3})\geq c.$
We remark that a constant could also be factored in front of $n^{1/3}$ inside the probability. For brevity, we do not carry this out throughout our proofs.
@@ 3 Bounds for the furthest distance a car can come from
First we divide the interval $(\varepsilon,\frac{\pi}{2}-\varepsilon)$ into disjoint sub-intervals. Let $c_{n}=\frac{n^{1/3}}{\lfloor n^{1/3}\rfloor}$. Then $\{c_{n}\}$ is a uniformly bounded sequence. Now, for $1\leq i\leq\lfloor n^{1/3}\rfloor$ we partition the interval $(\varepsilon,\frac{\pi}{2}-\varepsilon$) into intervals $A_{i}$ each of equal length $\frac{c_{n}(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$. For $1\leq i\leq\lfloor n^{1/3}\rfloor$ we define the following random variables.
$D_{i}:=\max\{m\in\mathbb{N}:\exists u\in\mathcal{L}_{-m}\text{ such that }\theta_{u}\in A_{i}\text{ and }\boldsymbol{0}\in\Gamma_{u}^{\theta_{u}}\}.$
@@
$D_{i}^{v}:=\max\{m\in\mathbb{N}:\exists u\in\mathcal{L}_{-n}\text{ such that }\theta_{u}\in A_{i}\text{ and }v\in\Gamma_{u}^{\theta_{u}}\}.$
Clearly by the translation invariance of the model, for all $v\in V$ we have
@@
Figure 3: To prove the upper bound in Theorem 2.5 we fix $1\leq i\leq\lfloor n^{1/3}\rfloor$. The event that there are more than $\ell$ many $u\in\mathcal{L}_{-n}$ with $\theta_{u}\in A_{i}$ and $\Gamma_{u}^{\theta_{u}}$ intersecting $V$ can happen in two ways. Either the above $u$ lies in $\mathcal{L}_{-n}\setminus\widetilde{V}$. For sufficiently large $\ell$ this will imply large transversal fluctuation for either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$ (the case that $\Gamma_{v_{1}}^{\alpha_{i}}$ can have large transversal fluctuation is shown in the figure. If for $u\in\mathcal{L}_{-n}\setminus\widetilde{V}$ as in the figure $,\Gamma_{u}^{\theta_{u}}$ intersects $V$ then by planarity of geodesics $\Gamma^{\alpha_{i}}_{v_{i}}$ will have large transversal fluctuation from the line $\mathcal{J}^{\alpha_{i}}_{v_{1}}$). Proposition 6.3 gives a stretched exponential upper bound for this event. The other possibility is there are more than $\ell$ distinct vertices inside on $V$ that carries a geodesic from $\widetilde{V}$. Proposition 6.4 gives a stretched exponential upper bound for this event.
@@
Let $\mathcal{J}_{v_{1}}^{\alpha_{i}}$ (resp. $\mathcal{J}_{v_{2}}^{\beta_{i}}$) be the straight line starting from $v_{1}$ (resp. $v_{2}$) in the direction $\alpha_{i}$ (resp. $\beta_{i}$). Now, by planarity the event $\mathcal{A}_{i}$ implies either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$ will have large traversal fluctuation on $\mathcal{L}_{0}$ from the lines $\mathcal{J}_{v_{1}}^{\alpha_{i}}$ and $\mathcal{J}_{v_{2}}^{\beta_{i}}$ respectively. (see Figure 3). Since, $(\beta_{i}-\alpha_{i})$ is of order $\frac{1}{n^{1/3}}$, for sufficiently large $\ell$ (depending on $\varepsilon$), the event $\mathcal{A}_{i}$ will imply that either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$ will have transversal fluctuation larger than $\frac{\ell^{1/96}n^{2/3}}{4}$. More precisely let $\mathcal{J}_{v_{1}}^{\alpha_{i}}(0)$ (resp. $\mathcal{J}_{v_{2}}^{\beta_{i}}(0)$) denote the intersection point of the corresponding lines with $\mathcal{L}_{0}$. Similarly let $\Gamma^{\alpha_{i}}_{v_{1}}(0)$ (resp. $\Gamma^{\beta_{i}}_{v_{2}}(0)$) denote the intersection points of the corresponding geodesics with $\mathcal{L}_{0}.$ Then
$\mathcal{A}_{i}\subset\left\{\Gamma^{\alpha_{i}}_{v_{1}}(0)-\mathcal{J}_{v_{1}}^{\alpha_{i}}(0)\geq\frac{\ell^{1/96}n^{2/3}}{4}\right\}\bigcup\left\{\mathcal{J}_{v_{2}}^{\beta_{i}}(0)-\Gamma^{\beta_{i}}_{v_{2}}(0)\geq\frac{\ell^{1/96}n^{2/3}}{4}\right\}.$
By Proposition 6.3 we have upper bounds for the events on the right side. For sufficiently large $\ell$ we have
@@
Now, we consider the event $\{\widehat{D_{i}}\geq\ell\}\cap\mathcal{A}_{i}^{c}$. On this event there are at least $\ell$ distinct vertices on $V$ that has a geodesic coming from $\widetilde{V}.$ Using Proposition 6.4 we have for sufficiently large $\ell$ and $n$
@@
So, finally we have that there exists $\ell_{0}$ (depending only on $\varepsilon$) and $\widetilde{C}$ (depending only on $\varepsilon$) such that for all large $n$
@@
So, from (3.2) we have that there exists $C>0$ (depending only on $\varepsilon$) such that for all $1\leq i\leq\lfloor n^{1/3}\rfloor$ we have
@@
We again consider disjoint sub-intervals of $(\varepsilon,\frac{\pi}{2}-\varepsilon)$. First we fix large constant $M$ which will be chosen later. For $1\leq i\leq[\frac{n^{1/3}}{2M}]$, we consider the intervals $A_{i}\subset(\varepsilon,\frac{\pi}{2}-\varepsilon)$, each of length $\frac{(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$ and midpoint $\frac{Mi(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$. Note that, unlike the previous case we are not partitioning the interval $(\varepsilon,\frac{\pi}{2}-\varepsilon$) this time. Instead we are taking points which are $\frac{M(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$ distance apart from each other and we are taking intervals of length $\frac{(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$ around them. We do this so that the constants that we get at the end depend only on $\varepsilon$ but not on $M$. This will be elaborated later. Corresponding to each $A_{i}$ we define the following random variables:
$D_{i}:=\max\left\{m\in\mathbb{N}:\exists u\in\mathcal{L}_{-n}\text{ such that }\theta_{u}\in A_{i}\text{ and }\boldsymbol{0}\in\Gamma_{u}^{\theta_{u}}\right\}.$
Note that $D_{i}$ is defined exactly in the same way as we defined in the proof of the upper bound. The only difference is that the definition of the intervals $A_{i}$ are different now. Clearly,
@@
So, applying inclusion-exclusion principle and considering the first two terms (see Proposition 6.1) we have
$\mathbb{P}(D\geq n)\geq\sum_{i=1}^{[\frac{n^{1/3}}{2M}]}\mathbb{P}\left(\left\{D_{i}\geq n\right\}\right)-\sum_{1\leq i<j\leq[\frac{n^{1/3}}{2M}]}\mathbb{P}\left(\left\{D_{i}\geq n\right\}\cap\left\{D_{j}\geq n\right\}\right).$ (3.4)
@@
$D_{i}^{v}:=\max\left\{m\in\mathbb{N}:\exists u\in\mathcal{L}_{-n}\text{ such that }\theta_{u}\in A_{i}\text{ and }v\in\Gamma_{u}^{\theta_{u}}\right\}.$
Figure 4: To prove the lower bound for the first sum in (3.4) we fix an $i$. Using Proposition 6.3 we can choose $M$ large enough so that on a positive probability event $\Gamma_{v_{1}^{\prime}}^{\beta_{i}}$ and $\Gamma_{v_{2}^{\prime}}^{\alpha_{i}}$ intersect $\mathcal{L}_{0}$ on $V_{M}$. Further, we consider an independent event on which there is at least one vertex on $V^{\prime}$ that chooses an angle in $A_{i}$. Due to central limit theorem the later event is a positive probability event. Hence, the intersection of these two events is a positive probability event and on this intersection $\widehat{D}_{i}\geq 1$.
@@
$\mathcal{C}_{i}:=\left\{\Gamma_{v_{1}^{\prime}}^{\beta_{i}}(0)\in V_{M}\text{ and }\Gamma_{v_{2}^{\prime}}^{\alpha_{i}}(0)\in V_{M}\right\}.$
Again as $(\beta_{i}-\alpha_{i})$ is of order $\frac{1}{n^{1/3}},$ the complement of the event $\mathcal{C}_{i}$ will imply that either $\Gamma_{v_{1}^{\prime}}^{\beta_{i}}$ or $\Gamma_{v_{2}^{\prime}}^{\alpha_{i}}$ will have transversal fluctuation more than $\frac{M}{2}n^{2/3}$. So, using Proposition 6.3 we can choose $M$ large enough so that
@@
$\mathcal{D}_{i}:=\{\exists u\in V^{\prime}\text{ such that }\theta_{u}\in A_{i}\}=\left\{\sum_{u\in V^{\prime}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}\geq 1\right\}.$
So, we consider the random variable $X_{i}:=\sum_{u\in V^{\prime}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}.$ We have $\mathbb{E}(X_{i})=\frac{\lceil n^{2/3}\rceil}{n^{1/3}}$ and $\operatorname{{\mathbb{V}ar}}(X_{i})=(\frac{1}{n^{1/3}}-\frac{1}{n^{2/3}})\lceil n^{2/3}\rceil.$ By the Berry-Esseen inequality (Proposition 6.2) we have for all $n$ there exists a constant $c>0$ such that
$\left|\mathbb{P}(X_{i}-\mathbb{E}(X_{i})\geq 0)-\frac{1}{2}\right|\leq\frac{c}{(n^{1/3}-1)^{3/2}n^{1/3}}.$
In particular, note that in this case the i.i.d. random variables in Proposition 6.2 are $\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}$ and we take $x$ in the statement of Proposition 6.2 to be $0$ Hence, for sufficiently large $n$ we have
$\mathbb{P}(X_{n}-\mathbb{E}(X_{n})\geq 0)\geq\frac{1}{4}.$
@@
Clearly, by planarity of geodesics (see Section 2 for definition), on the event $\mathcal{C}_{i}\cap\mathcal{D}_{i},\widehat{D_{i}}\geq 1.$ So, from (3.5) we have that there exists $c_{1}>0$ such that